7x+2y=10 -7x+y=-16
Answer by subudear(62)

7x+2y=10 ...(1)
-7x+y=-16 ...(2)
from equation (2) we get
y = -16 + 7x
replace value of y into equation (1) we get
7x + 2(-16 + 7x) = 10
7x - 32 + 14x = 10
21x = 10 + 32
x = 42/21
x = 2
replace value of x in equation (1) we get
7(2) + 2y = 10
2y = 10 -14
y = -4/2
y = -2
so solution for both the equations is x = 2, y = -2